OpenStudy (anonymous):

Translate to algebriac expression. The product of 29% and some number The translation is? (Type the percentage as a decimal. use b to represent some number.)

OpenStudy (anonymous):

I got b*0.29

OpenStudy (anonymous):

29% = 29/100 = 0.29 so if the number is b then product of 29% and b = 0.29 * b = 0.29b

OpenStudy (anonymous):

0.29b